OpenStudy (anonymous):

Find the standard form equation for a circle with center point at (6,-7) and diameter of 10

OpenStudy (akashdeepdeb):

Hello!! Welcome to OpenStudy!! In your question you are given the centre (x1,y1) = (6,-7) And the diamter as 10 or the radius as the half of that. Put these values in the standard equation of the circle ie. (x-x1)^2 + (y-y1)^2 = r^2 And you'll get your answer!! :D Understood? :)